Solve Problems Using Linear Systems
7. Tyler rows 10 km downstream in 2 h. On the return trip, it takes hom 4 h to travel 8 km. Determine his average rowing speed and the speed of the current.
(answrs are: average rowing speed 3.5 km/h, speed of current 1.5 km/h, but i dont know how to solve it?)

Answer by stanbon(75887)

Tyler rows 10 km downstream in 2 h. On the return trip, it takes hom 4 h to travel 8 km. Determine his average rowing speed and the speed of the current.
------------------------------------------
Downstream DATA:
distance = 10 km ; time = 2h ; rate = 5 km/h
---------------------------------
Upstream DATA:
distance = 8 km ; time = 4 hr ; rate = 2 km/h
--------------------------------
Equations:
b + c = 5
b - c = 2
------------
Add to solve for "b":
2b = 7
b = 3.5 km/h (speed of the boat in still water)
-------------------
substitute into b+c = 5 to solve for "c":
3.5 + c = 5
c = 1.5 km/h speed of the current)
============================================
cheers,
Stan H.